OpenStudy (anonymous):

Solve the equation or formula for the indicated variable. s=6r^5t, for t

OpenStudy (anonymous):

t = S/5r^2

OpenStudy (anonymous):

Can you show the work I'm so confused.

OpenStudy (anonymous):

Well you want to isolate t, that means you want everything except t to be on one side of the equals-sign. The method for that is to do the reverse operations wherever t is not alone, so in this t is multiplied by 5r2 so we divide both sides by 5r2, and the result is S5r2=5r2t5r2 On the right side we can cancel 5r2, so the answer is t=S5r2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome! be sure to ask if you need help with this cool stuff again!

OpenStudy (anonymous):

I'll probably have a few more questions.

OpenStudy (anonymous):

ok

OpenStudy (phi):

your equation is not clear. Is it \[ s=6r^5t \] or \[ s=6r^{5t}\]

OpenStudy (anonymous):

@phi the first one...

OpenStudy (phi):

ok, then to solve for t, divide both sides by 6 r^5 \[ s=6r^5t \\ \frac{s}{6r^5} = \frac{\cancel{6r^5}}{\cancel{6r^5}}t \] you get \[ t = \frac{s}{6r^5} \] which can be written in different ways. example \[ t = \frac{1}{6} s r^{-5} \]

OpenStudy (anonymous):

i already did